OpenStudy (pulsified333):

Two fair die are rolled. (1) What is the probability they sum to 10, 7, or 4? (2) What is the probability they sum to at least 11?

OpenStudy (anonymous):

So for two dices, there are 6*6 = 36 possible combinations right?

OpenStudy (pulsified333):

yes

OpenStudy (anonymous):

how many combinations add up to 4?

OpenStudy (pulsified333):

2

OpenStudy (anonymous):

1,3 3,1 2,2 there are 3

OpenStudy (pulsified333):

okay

OpenStudy (anonymous):

so for 4, it is 3/36 = 1/12

OpenStudy (pulsified333):

correct

OpenStudy (anonymous):

so you can do the same for 7 and 10 to for (1).

OpenStudy (pulsified333):

are you saying that there are 6 ways to add up to 6?

OpenStudy (anonymous):

yes, so its probability is 1/6

OpenStudy (pulsified333):

btw how did you get 1/12 for 4?

OpenStudy (anonymous):

sorry sorry no there's 5 ways to add up to 6

OpenStudy (anonymous):

1,5 2,4 3,3 4,2 5,1

OpenStudy (pulsified333):

The number is seven my bad

OpenStudy (anonymous):

Oh yes

OpenStudy (pulsified333):

can you explain how you got 3/36 though for 4?

OpenStudy (anonymous):

3 combinations to get 4 / all the possible combinations = 3/36.

OpenStudy (pulsified333):

how is it 36 though

OpenStudy (anonymous):

there are 6 possibilities on the first dice & 6 possibilities on the second dice. 6 * 6 = 36

OpenStudy (pulsified333):

oh never mind. sorry I didn't think about that, I've had a really long day. You are right

OpenStudy (anonymous):

oh ok! :D

OpenStudy (pulsified333):

so in total it should be 6/36+3/36+3/36=12/36=1/3

OpenStudy (anonymous):

and for (2), you just have to find the probability that the sum is 12 and subtract it from 1

OpenStudy (anonymous):

and yes

OpenStudy (pulsified333):

why subtract one for the sum is 12?

OpenStudy (anonymous):

at least 11 is the same as not 12, becuase the biggest sum is 12

OpenStudy (pulsified333):

right. So that means 11 and 12?

OpenStudy (anonymous):

so, "at least 11" refers to sums 2-11. so the only sum that is not considered is 12.

OpenStudy (pulsified333):

oh

OpenStudy (anonymous):

do you get that?

OpenStudy (anonymous):

so what you have to do is 1 - (probability of getting 12) = probability of all sums at least 11

OpenStudy (pulsified333):

so is that mean the answer is 35/36?

OpenStudy (anonymous):

yes :D

OpenStudy (pulsified333):

it says 35/36 is wrong but the first answer is right

OpenStudy (anonymous):

really? :,o let me think I'm sorry...

OpenStudy (pulsified333):

its okay

OpenStudy (anonymous):

AHHH I READ THE PROBLEM WRONG... YOU WERE RIGHT IT IS PROBABILITY OF 11 AND 12

OpenStudy (pulsified333):

so its 1/12

OpenStudy (anonymous):

I CAN'T BE MORE SORRY THAN THIS SORRY

OpenStudy (anonymous):

YES

OpenStudy (pulsified333):

dude its alright

OpenStudy (pulsified333):

That was right

OpenStudy (pulsified333):

:D thank you very much

OpenStudy (anonymous):

thank you :)